# Constrained Error-Correcting Codes for Efficient DNA Synthesis

### v1

@@ I Introduction
In the DNA synthesis process, a large number of DNA strands are commonly created in parallel [6, 7, 8]. Generally, the synthesis machine follows a fixed template supersequence $S$ and adds one nucleotide at a time to form the specific DNA strands. The synthesis process continues until the entire supersequence is exhausted. Consequently, the length of the fixed supersequence dictates the number of cycles required to generate all the strands, corresponding to the total synthesis time. This leads to an interesting constrained coding problem, first introduced by Lenz et al. [5] and subsequently studied in [9, 10, 11, 12, 13, 14, 15]. In [5], the authors demonstrated that the supersequences that maximize the information rate are the alternating quaternary sequences. Therefore, in this paper, we employ the alternating sequence $S=ATCGATCG\cdots$ as the fixed template supersequence for DNA synthesis. Moreover, DNA synthesis is prone to various types of errors, such as insertions, deletions, and substitutions. To mitigate those errors, two constraints are typically employed in practice. One is the $\ell$-runlength limit ($\ell$-RLL) [16, 17], which limits the length of consecutive identical nucleotides in a DNA sequence to less than $\ell$, where $\ell$ is a positive integer. The other is the $\epsilon$-balance [17, 18], which ensures that the ratio of $G$ and $C$ in a DNA sequence lies within the range $[0.5-\epsilon,0.5+\epsilon]$, where $\epsilon\in[0,0.5]$. To effectively reduce or correct errors in DNA synthesis, it is critical to take into account the constructions of constrained codes that satisfy $\ell$-RLL and/or $\epsilon$-balanced constraints, as well as constrained error-correcting codes (ECCs) capable of correcting errors while adhering to these two constraints.
TABLE I: Comparison of Different Constrained ECCs for DNA synthesis, where $r_{*},*\in\{\ell,T,(\ell,T),(\ell,\epsilon,T)\}$ denotes the redundancy of enumeration methods
Synthesis time $T$ Redundancy Computation Complexity Property Ref.
@@
Building upon advancements in the DNA synthesis machine, various studies have explored both the theoretical aspects and code constructions for DNA synthesis [9, 10, 11, 12, 5, 13, 14, 15]. Specifically, Sini et al. investigated optimizing DNA synthesis by using shortmers instead of single nucleotides in [9]. Makarychev et al. [10] studied batch optimization techniques to reduce the cost of large-scale DNA synthesis. Building on this work, Elishco et al. [11] advanced the theoretical understanding of DNA synthesis by modeling the process as a constrained system and analyzing the $\ell$-RLL constraint. Immink et al. [12] explored the trade-offs between several components required to synthesize multiple DNA strands in parallel.
We can observe that the synthesis time of an arbitrary DNA sequence with length $n$ is at most $T\leq 4n$ during DNA synthesis. Recent developments have demonstrated that introducing redundancy can significantly reduce the synthesis time. For instance, Lenz et al. [5] proposed an explicit construction that limits the synthesis time to $T\leq 2.5n$ for a DNA sequence with length $n$, using only one bit of redundancy. Chrisnata et al. [13] further enhanced the process by designing systematic ECCs that can correct single insertion/deletion (indel) error, while maintaining the same synthesis time of $T\leq 2.5n$, but with $\lceil\log_{4}n\rceil+5$ symbols of redundancy. Recently, Nguyen et al. [14] developed constrained ECCs that achieved a synthesis time of $T\leq 2.5n-1$ while ensuring the $\ell$-RLL constraint. The results from [15] demonstrated two families of codes capable of correcting synthesis defects. Additionally, many other studies have focused on constrained codes with $\ell$-RLL or/and $\epsilon$-balance for DNA storage systems [19, 20, 21, 22, 23]. However, these works have predominantly concentrated on DNA storage systems as a whole, rather than the DNA synthesis. Thus, the challenge of constructing constrained ECCs that simultaneously satisfy both the $\ell$-RLL and $\epsilon$-balance constraints ($(\ell,\epsilon)$-constraints) remains an open problem in the context of DNA synthesis.
To address this issue, we propose explicit constrained codes that incorporate both $\ell$-RLL and $\epsilon$-balance for DNA synthesis, with a fixed synthesis time $T$ satisfying $n\leq T\leq 4n$. These codes encode the binary information into the DNA sequences with polynomial-time encoding and decoding algorithms. By enumerating all valid sequences, the rate of our codes achieves the theoretical capacity. Furthermore, we develop constrained ECCs for DNA synthesis, which can correct one indel in DNA strands. We compare our constructions with some existing results in Table I.
@@ II Preliminaries
Let $\Sigma_{4}=\{1,2,3,4\}$. For a DNA system, we map the four different bases as follows: $A\rightarrow 1,T\rightarrow 2,C\rightarrow 3$ and $G\rightarrow 4$, and use the shifted modulo operator $(a\bmod 4)\in\{1,2,3,4\}$ throughout this paper. Let $\{a\}^{t}$ be a length-$t$ sequence formed by $a\in\Sigma_{4}$. For a DNA sequence $\mathbf{c}\in\Sigma_{4}^{n}$, we write
@@
where $a_{i}\neq a_{i+1},i\in[1,k),\sum_{i=1}^{k}t_{i}=n$ and ||||| | is a concatenation of two sequences (e.g., $\{a_{1}\}||\{a_{2}\}^{2}||\{a_{3}\}=a_{1}a_{2}a_{2}a_{3}$). Particularly, $\mathbf{c}||\emptyset=\mathbf{c}$. Each sub-sequence $\{a_{i}\}^{t_{i}}$ is called a run in $\mathbf{c}$, and $\{a_{k}\}^{t_{k}}$ is the last run. We say that $\mathbf{c}$ satisfies $\ell$-RLL if $t_{i}\in[1,\ell]$, $\forall~{}i\in[1,k]$. In this paper, we consider the set of length-$n$ DNA sequences with $\ell$-RLL, thus, all valid runs are included in the following set,
@@
For a run $\mathbf{s}=\{a\}^{t}\in\mathcal{S}$, denote the element and the length of $\mathbf{s}$ by $e(\mathbf{s})=a$ and $|\mathbf{s}|=t$, respectively. For a sequence $\mathbf{c}=(c_{1},c_{2},\dots,c_{n})\in\Sigma_{4}^{n}$, denote the GC-content and synthesis time of $\mathbf{c}$ by $G(\mathbf{c})$ and $T(\mathbf{c})$, respectively, where
@@
therein $\mathbb{1}_{\{*\}}$ equals $1$ when condition * is true, otherwise 00. We say that $\mathbf{c}$ satisfies $\epsilon$-balance if
$G(\mathbf{c})\in\mathcal{M}=[\lceil{(0.5-\epsilon)n}\rceil,\lfloor{(0.5+\epsilon)n}\rfloor].$
Particularly, when $G(\mathbf{c})=0.5n$, i.e., $\epsilon=0$, we say that $\mathbf{c}$ is balanced.
@@ III Constrained Codes for DNA synthesis
In this section, we construct capacity-achieving $\ell$-RLL and $(\ell,\epsilon)$-constrained codes for DNA synthesis, denoted by $\widetilde{\mathcal{C}}(n,\ell,T)$ and $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$, respectively, where $n$ and $T$ are the code length and synthesis time. We begin by partitioning all DNA sequences that satisfy the $\ell$-RLL constraint into some subsets, which are subsequently ordered according to specific rules. Following this, we develop ranking and unranking methods to facilitate the encoding and decoding processes, respectively. By enumerating all valid sequences, we ensure that the encoding and decoding algorithms of our codes have polynomial-time computational complexities.
@@ III-A $\ell$-RLL constrained codes for DNA synthesis
For a fixed $n\leq T\leq 4n$, denote the set of length-$n$ sequences satisfying $\ell$-RLL, synthesis time at most $T$ by $\widetilde{\mathcal{C}}(n,\ell,T)$. Denote all DNA sequences with length $n$, last run $\mathbf{s}$ and synthesis time at most $T$ by $\mathcal{C}(n,\mathbf{s},T)$. Thus, we can write
@@
$\mathcal{C}(n,\mathbf{s},T)=\bigcup_{\mathbf{s}^{\prime}\in\mathcal{S}:e(\mathbf{s})\neq e(\mathbf{s}^{\prime})}\mathcal{C}(n-|\mathbf{s}|,\mathbf{s}^{\prime},T^{\prime})||\mathbf{s},$
@@
For any two DNA sequences $\mathbf{c}_{1}\in\mathcal{C}(n,\mathbf{s}_{1},T_{1})$ and $\mathbf{c}_{2}\in\mathcal{C}(n,\mathbf{s}_{2},T_{2})$, where $\mathbf{s}_{1},\mathbf{s}_{2}\in\mathcal{S}$, we order the sequences logically by the length and element of the last run. Define $\mathbf{c}_{1}<\mathbf{c}_{2}$ if any of the following conditions are satisfied.
@@
Algorithm 1 $\mathrm{Enc}_{\ell,T}$: Unrank the $M$-th sequence in $\widetilde{\mathcal{C}}(n,\ell,T)$
@@
Algorithm 2 $\mathrm{Dec}_{\ell,T}$: Rank $\mathbf{c}$ in $\widetilde{\mathcal{C}}(n,\ell,T)$
@@
For given $n,\ell,\epsilon$ and a fixed $n\leq T\leq 4n$, the code $\widetilde{\mathcal{C}}(n,\ell,T)$ encodes a binary sequence of length $\lceil\log_{2}{N(n,\ell,T)}\rceil$ to a DNA sequence of length $n$ with $\ell$-RLL and synthesis time at most $T$. This results in a redundancy of $r_{\ell,T}=n-\lceil\log_{4}{N(n,\ell,T)}\rceil$. In addition, the encoding and decoding algorithms run in $O(n^{3}\ell^{2})$ time, thus, $\widetilde{\mathcal{C}}(n,\ell,T)$ is a capacity-achieving constrained code for DNA synthesis with polynomial-time complexity.
@@
The computational complexity includes the following two parts: i). Precomputing and storing of $N$ (offline preprocessing): We need to precompute and store $N(n^{\prime},\mathbf{s}^{\prime},T^{\prime})$ for all $n^{\prime}\in[1,n],\mathbf{s}^{\prime}\in\mathcal{S},T^{\prime}\in[1,T]$, i.e., there are $O(n^{2}\ell)$ states for $N$. Each $N(n^{\prime},\mathbf{s}^{\prime},T^{\prime})$ is a denary number of length $O(n)$. According to (1), the computation of $N(n,\mathbf{s},T)$ is based on the other $O(\ell)$ states, indicating the computational complexity for computing $N$ is $O(n^{3}\ell^{2})$. ii). Computation in Algorithms 1 and 2 for a given $N$ (online processing): Both Algorithms 1 and 2 iterate for at most $n$ times. In each iteration, the computation of $M_{1}$ dominates the complexity. As can be seen from Line 7 of Algorithm 1 and Line 8 of Algorithm 2, $\mathcal{S}$ is of size $O(\ell)$ for each iteration, and $N(n,\mathbf{s},T)$ is a denary number of length $O(n)$. Therefore, the computation of $M_{1}$ has computational complexity $O(n^{2}\ell)$. ∎
@@ III-B $(\ell,\epsilon)$-constrained codes for DNA synthesis
For a fixed $n\leq T\leq 4n$, denote the set of length-$n$ sequences satisfying $(\ell,\epsilon)$-constraints, synthesis time at most $T$ by $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$. Denote all sequences with length $n$, last run $\mathbf{s}$, GC-content $m$, synthesis time at most $T$ by $\mathcal{C}(n,\mathbf{s},m,T)$. Thus, we can write
$\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)=\bigcup_{\mathbf{s}\in\mathcal{S},m\in\mathcal{M}}\mathcal{C}(n,\mathbf{s},m,T).$
@@
$\mathcal{C}(n,\mathbf{s},m,T)=\bigcup_{\mathbf{s}^{\prime}\in\mathcal{S}:e(\mathbf{s})\neq e(\mathbf{s}^{\prime})}\mathcal{C}(n-|\mathbf{s}|,\mathbf{s}^{\prime},m-G(\mathbf{s}),T^{\prime})||\mathbf{s},$
@@
To order all the sequences in $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$, we order the sequences logically in each subset $\mathcal{C}(n,\mathbf{s},m,T)$ by GC-content first and then the length and element of the last run. For any two sequences $\mathbf{c}_{1}\in\mathcal{C}(n,\mathbf{s}_{1},m_{1},T_{1})$ and $\mathbf{c}_{2}\in\mathcal{C}(n,\mathbf{s}_{2},m_{2},T_{2})$, define $\mathbf{c}_{1}<\mathbf{c}_{2}$ if any of the following conditions are satisfied.
@@
In the ranking method, we determine the ordering of a given DNA sequence. This implies that there exists a unique integer $M\in[1,N(n,\ell,\epsilon,T)]$ corresponding to $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$. We define the function $\mathrm{rank}(\mathbf{c},n,\mathcal{S}\times\mathcal{M},T)$ as outlined in Algorithm 2, which enables us to compute $M_{1}$ in Line 8. Therein, $M_{1}$ is defined in (8), and $\mathbf{s},m$ are the last run and GC-content of $\mathbf{c}$, respectively. Then we only need to return $\mathrm{rank}(P_{n-|\mathbf{s}|}(\mathbf{c}),n-|\mathbf{s}|,\mathcal{S}^{\prime}\times\{m^{\prime}\},T^{\prime})+M_{1}$.
@@
For given $n,\ell,\epsilon$ and a fixed $n\leq T\leq 4n$, the code $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$ encodes a binary sequence of length $\lceil\log_{2}{N(n,\ell,\epsilon,T)}\rceil$ to a DNA sequence of length $n$ with $\ell$-RLL, $\epsilon$-balance and synthesis time at most $T$. This results in a redundancy of $r_{\ell,\epsilon,T}=n-\lceil\log_{4}{N(n,\ell,\epsilon,T)}\rceil$. In addition, the encoding and decoding algorithms run in $O(n^{4}\ell^{2})$ time, thus $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$ is a capacity-achieving constrained code for DNA synthesis with polynomial-time complexity.

### v3

@@ I Introduction
In the DNA synthesis process, a large number of DNA strands are commonly created in parallel [6, 7, 8]. Generally, the synthesis machine follows a fixed template supersequence $S$ and adds one nucleotide at a time to form the specific DNA strands. The synthesis process continues until the entire supersequence is exhausted. Consequently, the length of the fixed supersequence dictates the number of cycles required to generate all the strands, corresponding to the total synthesis time. This leads to an interesting constrained coding problem, first introduced by Lenz et al. [5] and subsequently studied in [9, 10, 11, 12, 13, 14, 15]. In [5], the authors demonstrate that the supersequences that maximize the information rate are the alternating quaternary sequences. Therefore, in this paper, we employ the alternating sequence $S=ATCGATCG\cdots$ as the fixed template supersequence for DNA synthesis. Moreover, DNA synthesis is prone to various types of errors, such as insertions, deletions, and substitutions. To mitigate those errors, two constraints are typically employed in practice. One is the $\ell$-runlength limit ($\ell$-RLL) [16, 17], which restricts the length of consecutive identical nucleotides in a DNA sequence to be no greater than $\ell$, where $\ell$ is a positive integer. The other is the $\epsilon$-balance [17, 18], which ensures that the ratio of $G$ and $C$ in a DNA sequence lies within the range $[0.5-\epsilon,0.5+\epsilon]$, where $\epsilon\in[0,0.5]$. To effectively reduce or correct errors in DNA synthesis, it is critical to take into account the constructions of constrained codes that satisfy $\ell$-RLL and/or $\epsilon$-balanced constraints, as well as constrained error-correcting codes (ECCs) capable of correcting errors while adhering to these two constraints.
TABLE I: Comparison of different constrained ECCs for DNA synthesis, where $r_{*},*\in\{\ell,T,(\ell,T),(\ell,\epsilon,T)\}$ denotes the redundancy of enumeration methods.
Synthesis time $T$ Redundancy Computational Complexity Property Ref.
@@
Building upon the advancements in the DNA synthesis machine, various studies have explored both the theoretical aspects and code constructions for DNA synthesis [9, 10, 11, 12, 13, 14, 15]. Specifically, Sini et al. [9] investigated optimizing DNA synthesis by using shortmers instead of single nucleotides. Makarychev et al. [10] studied batch optimization techniques to reduce the cost of large-scale DNA synthesis. Building on this work, Elishco et al. [11] advanced the theoretical understanding of DNA synthesis by modeling the process as a constrained system and analyzing the $\ell$-RLL constraint. Immink et al. [12] explored the trade-offs between several components required to synthesize multiple DNA strands in parallel.
We can observe that the synthesis time of an arbitrary DNA sequence with length $n$ is at most $T\leq 4n$ during DNA synthesis. Recent developments have demonstrated that introducing redundancy can significantly reduce the synthesis time. For instance, Lenz et al. [5] proposed an explicit construction that limits the synthesis time to $T\leq 2.5n$ for a DNA sequence with length $n$, using only one bit of redundancy. Chrisnata et al. [13] further enhanced the process by designing systematic ECCs that can correct single insertion/deletion (indel) error, while maintaining the same synthesis time of $T\leq 2.5n$, but with $\lceil\log_{4}n\rceil+5$ symbols of redundancy. Recently, Nguyen et al. [14] have developed constrained ECCs that achieve a synthesis time of $T\leq 2.5n-1$ while ensuring the $\ell$-RLL constraint. The results from [15] demonstrate two families of codes capable of correcting synthesis defects. Additionally, many other studies have focused on constrained codes with $\ell$-RLL or/and $\epsilon$-balance for DNA storage systems [19, 20, 21, 22, 23]. However, these works have predominantly concentrated on DNA storage systems as a whole, rather than DNA synthesis. Thus, the challenge of constructing constrained ECCs that simultaneously satisfy both the $\ell$-RLL and $\epsilon$-balance constraints ($(\ell,\epsilon)$-constraints) remains an open problem in the context of DNA synthesis.
To address this issue, we propose explicit constrained codes that incorporate both $\ell$-RLL and $\epsilon$-balance for DNA synthesis, with a fixed synthesis time $T$ satisfying $n\leq T\leq 4n$. These codes encode the binary information into the DNA sequences with polynomial-time encoding and decoding algorithms. By enumerating all valid sequences, the rate of our codes achieves the theoretical capacity. Furthermore, we develop constrained ECCs for DNA synthesis, which can correct one indel in the DNA strand. We compare the designed constructions with some existing results in Table I.
@@ II Preliminaries
Let $\Sigma_{4}=\{1,2,3,4\}$. For a DNA system, we map the four different bases as follows: $A\rightarrow 1,T\rightarrow 2,C\rightarrow 3$ and $G\rightarrow 4$, and use the shifted modulo operator $(a\bmod 4)\in\{1,2,3,4\}$ throughout this paper. Denote $[a,b]=\{a,a+1,\dots,b\}$ and $[a,b)=\{a,a+1,\dots,b-1\}$ as two order sets for any two non-negative integers $a,b$ with $a<b$, respectively. Let $\{a\}^{t}$ be a length-$t$ sequence formed by $a\in\Sigma_{4}$. For a DNA sequence $\mathbf{c}\in\Sigma_{4}^{n}$, we write
@@
where $a_{i}\neq a_{i+1},i\in[1,k),\sum_{i=1}^{k}t_{i}=n$ and ||||| | denotes a concatenation of two sequences (e.g., $\{a_{1}\}||\{a_{2}\}^{2}||\{a_{3}\}=a_{1}a_{2}a_{2}a_{3}$). Particularly, $\mathbf{c}||\emptyset=\mathbf{c}$. Each sub-sequence $\{a_{i}\}^{t_{i}}$ is called a run in $\mathbf{c}$, and $\{a_{k}\}^{t_{k}}$ is the last run. We say that $\mathbf{c}$ satisfies $\ell$-RLL if $t_{i}\in[1,\ell]$, $\forall~{}i\in[1,k]$. In this paper, we consider the set of length-$n$ DNA sequences with $\ell$-RLL, thus, all valid runs are included in the following set,
@@
For a run $\mathbf{s}=\{a\}^{t}\in\mathcal{S}$, denote the symbol and the length of $\mathbf{s}$ by $e(\mathbf{s})=a$ and $|\mathbf{s}|=t$, respectively. For a sequence $\mathbf{c}=(c_{1},c_{2},\dots,c_{n})\in\Sigma_{4}^{n}$, denote the GC-content and synthesis time of $\mathbf{c}$ by $G(\mathbf{c})$ and $T(\mathbf{c})$, respectively, where
@@
therein $\mathbb{1}_{\{*\}}$ equals $1$ when condition * is true, otherwise it is 00. We say that $\mathbf{c}$ satisfies $\epsilon$-balance if
$G(\mathbf{c})\in\mathcal{M}=[\lceil{(0.5-\epsilon)n}\rceil,\lfloor{(0.5+\epsilon)n}\rfloor].$ (2)
Particularly, when $n$ is even and $G(\mathbf{c})=0.5n$, i.e., $\epsilon=0$, we say that $\mathbf{c}$ is balanced.
@@ III Constrained Codes for DNA Synthesis
In this section, we construct capacity-achieving $\ell$-RLL and $(\ell,\epsilon)$-constrained codes for DNA synthesis, denoted by $\widetilde{\mathcal{C}}(n,\ell,T)$ and $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$, respectively, where $n$ and $T$ are the code length and synthesis time. We begin by partitioning the set of DNA sequences that satisfy the $\ell$-RLL constraint into some subsets, which are subsequently ordered according to specific rules. Following this, we develop unranking and ranking methods to facilitate the encoding and decoding processes, respectively. Both the encoding and decoding algorithms of the developed codes have polynomial-time computational complexities.
@@ III-A $\ell$-RLL Constrained Codes for DNA Synthesis
For a fixed $T$ with $n\leq T\leq 4n$, denote the set of length-$n$ sequences satisfying the $\ell$-RLL constraint with synthesis time at most $T$ by $\widetilde{\mathcal{C}}(n,\ell,T)$. Denote all the DNA sequences with length $n$, last run $\mathbf{s}$ and synthesis time at most $T$ by $\mathcal{C}(n,\mathbf{s},T)$. Thus, we can write
@@
$\mathcal{C}(n,\mathbf{s},T)=\{\mathbf{c}^{\prime}||\mathbf{s}:\mathbf{c}^{\prime}\in\mathcal{C}(n-|\mathbf{s}|,\mathbf{s}^{\prime},T^{\prime}),\mathbf{s}^{\prime}\in\mathcal{S}$
 $\text{~{}with~{}}e(\mathbf{s}^{\prime})\neq e(\mathbf{s})\},$
@@
For any pairs of DNA sequences $\mathbf{c}_{1}\in\mathcal{C}(n,\mathbf{s}_{1},T_{1})$ and $\mathbf{c}_{2}\in\mathcal{C}(n,\mathbf{s}_{2},T_{2})$, where $\mathbf{s}_{1},\mathbf{s}_{2}\in\mathcal{S}$, we order the sequences logically by the length and symbol of the last run. Define $\mathbf{c}_{1}<\mathbf{c}_{2}$ if any of the following conditions are satisfied.
@@
Algorithm 1 $\mathrm{Enc}_{\ell,T}$: Unrank the $M$-th sequence in $\widetilde{\mathcal{C}}(n,\ell,T)$.
@@
Algorithm 2 $\mathrm{Dec}_{\ell,T}$: Rank $\mathbf{c}$ in $\widetilde{\mathcal{C}}(n,\ell,T)$.
@@
For given $n,\ell,\epsilon$ and a fixed $T$ with $n\leq T\leq 4n$, the code $\widetilde{\mathcal{C}}(n,\ell,T)$ encodes a binary sequence of length $\lceil\log_{2}{N(n,\ell,T)}\rceil$ to a DNA sequence of length $n$ with $\ell$-RLL and synthesis time at most $T$. This results in a redundancy of $r_{\ell,T}=n-\lceil\log_{4}{N(n,\ell,T)}\rceil$. In addition, the encoding and decoding algorithms run in $O(n^{3}\ell^{2})$ time, thus, $\widetilde{\mathcal{C}}(n,\ell,T)$ is a capacity-achieving constrained code for DNA synthesis with polynomial-time complexity.
@@
The computational complexity includes the following two parts: i). Precomputing and storing of $N$ (offline preprocessing): We need to precompute and store $N(n^{\prime},\mathbf{s}^{\prime},T^{\prime})$ for all $n^{\prime}\in[1,n],\mathbf{s}^{\prime}\in\mathcal{S},T^{\prime}\in[1,T]$, i.e., there are $O(n^{2}\ell)$ states for $N$. Each $N(n^{\prime},\mathbf{s}^{\prime},T^{\prime})$ is a denary number of length $O(n)$. According to (1), the computation of $N(n,\mathbf{s},T)$ is based on the other $O(\ell)$ states, indicating the computational complexity for computing $N$ is $O(n^{3}\ell^{2})$. ii). Computation in Algorithms 1 and 2 for a given $N$ (online processing): Both Algorithms 1 and 2 iterate at most $n$ times. In each iteration, the computation of $M_{1}$ dominates the complexity. As can be seen from Line 7 of Algorithm 1 and Line 8 of Algorithm 2, the set $\mathcal{S}$ is of size $O(\ell)$ for each iteration, and $N(n,\mathbf{s},T)$ is a denary number of length $O(n)$. Therefore, the computation of $M_{1}$ has computational complexity $O(n^{2}\ell)$. ∎
@@ III-B $(\ell,\epsilon)$-Constrained Codes for DNA Synthesis
For a fixed $T$ with $n\leq T\leq 4n$, denote the set of length-$n$ sequences satisfying the $(\ell,\epsilon)$-constraints with synthesis time at most $T$ by $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$. Denote all the sequences with length $n$, last run $\mathbf{s}$, GC-content $m$, and synthesis time at most $T$ by $\mathcal{C}(n,\mathbf{s},m,T)$. Thus, we can write
$\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)=\bigcup_{\mathbf{s}\in\mathcal{S},m\in\mathcal{M}}\mathcal{C}(n,\mathbf{s},m,T),$
where $\mathcal{S}$ and $\mathcal{M}$ are defined in (1) and (2), respectively.
@@
$\mathcal{C}(n,\mathbf{s},m,T)=\{\mathbf{c}^{\prime}||\mathbf{s}:\mathbf{c}^{\prime}\in\mathcal{C}(n-|\mathbf{s}|,\mathbf{s}^{\prime},m-G(\mathbf{s}),T^{\prime}),$
$\mathbf{s}^{\prime}\in\mathcal{S}\text{~{}with~{}}e(\mathbf{s}^{\prime})\neq e(\mathbf{s})\},$
@@
To order all the sequences in $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$, we order the sequences logically in each subset $\mathcal{C}(n,\mathbf{s},m,T)$ by GC-content first, then the length and symbol of the last run. For any pairs of sequences $\mathbf{c}_{1}\in\mathcal{C}(n,\mathbf{s}_{1},m_{1},T_{1})$ and $\mathbf{c}_{2}\in\mathcal{C}(n,\mathbf{s}_{2},m_{2},T_{2})$, define $\mathbf{c}_{1}<\mathbf{c}_{2}$ if any of the following conditions are satisfied.
@@
In the ranking method, we determine the ordering of a given DNA sequence. This implies that there exists a unique integer $M\in[1,N(n,\ell,\epsilon,T)]$ corresponding to a DNA sequence $\mathbf{c}\in\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$. We define the function $\mathrm{rank}(\mathbf{c},n,\mathcal{S}\times\mathcal{M},T)$ as outlined in Algorithm 2, which enables us to compute $M_{1}$ in Line 8. Therein, $M_{1}$ is defined in (9), and $\mathbf{s},m$ are the last run and GC-content of $\mathbf{c}$, respectively. Then we only need to return $\mathrm{rank}(P_{n-|\mathbf{s}|}(\mathbf{c}),n-|\mathbf{s}|,\mathcal{S}^{\prime}\times\{m^{\prime}\},T^{\prime})+M_{1}$, where $m^{\prime}=m-G(\mathbf{s})$.
@@
For given $n,\ell,\epsilon$ and a fixed $T$ with $n\leq T\leq 4n$, the code $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$ encodes a binary sequence of length $\lceil\log_{2}{N(n,\ell,\epsilon,T)}\rceil$ to a DNA sequence of length $n$ with $\ell$-RLL, $\epsilon$-balance and synthesis time at most $T$. This results in a redundancy of $r_{\ell,\epsilon,T}=n-\lceil\log_{4}{N(n,\ell,\epsilon,T)}\rceil$. In addition, the encoding and decoding algorithms run in $O(n^{4}\ell^{2})$ time, thus $\widetilde{\mathcal{C}}(n,\ell,\epsilon,T)$ is a capacity-achieving constrained code for DNA synthesis with polynomial-time complexity.